### v3

# Vision Language Models See What You Want but not What You See

Abstract
Knowing others’ intentions and taking others’ perspectives are two core components of human intelligence typically considered as instantiations of theory of mind. Infiltrating machines with these abilities is an important step towards building human-level artificial intelligence. We here investigate intentionality understanding and perspective-taking in Vision Language Models and, for the purpose, we have created IntentBench and PerspectBench datasets, which contain over 400 cognitive experiments grounded in real-world scenarios and classic cognitive tasks. Surprisingly, we find that VLMs achieve high performance in intentionality understanding but lower performance in perspective-taking using our two datasets. This challenges the common belief in the cognitive science literature that perspective-taking at the corresponding modality is necessary for intentionality understanding.
1 Introduction
Intentionality is the capacity of the mind to be directed toward, represent, or stand for objects, properties, or states of affairs for further executable actions [Anscombe, 1956]. To say one has intentionality is to say that one has a mental representation of content for action [Dennett, 1969, Kripke, 1982]. To say one could understand intentionality is to say one has the capacity to comprehend the mental representation of content for action in another mind [Premack and Woodruff, 1978, Rosenthal, 1991]. This capacity has been seen as a key distinction between humans and machines [Searle, 1980], as proposed in the Chinese Room Argument. In this thought experiment, a person inside a room follows a dictionary to translate English into Chinese without understanding the intentional meaning of the symbols, illustrating how machines can process information syntactically without true comprehension. It is argued that despite well manipulation of language symbols, machine still lacks true understanding or consciousness because it cannot grasp the intentions behind the words, the mental content for action conveyed by symbols [Sellars, 1956, Grice, 1957, Davidson, 1987]. To truly understand intentional meaning, theory-of-mind, the ability to simulate the mental content of others, is required [Premack and Woodruff, 1978, Dennett, 1987]. From there, it is concluded that it is impossible to build a machine without theory-of-mind yet still having the capacity of intentionality understanding [Searle, 1980].
Figure 1: VLMs Performance on IntentBench and PerspectBench as Compared to Human Baselines
Theory-of-mind is commonly understood to be grounded in perspective-taking, the ability to cognitively undertake the perspective of another [Barnes-Holmes et al., 2004]. In particular, level-1 perspective-taking refers to the acknowledgment that different people can see different things, which is the most primitive form of theory-of-mind. Building on such level-1 knowledge is level-2 perspective-taking, which is the ability to know that people can see the same things differently. The acquisition of level-2 perspective-taking, referred to as the concept of perspectives in Piagetian developmental psychology, marks a milestone of human cognitive development through elimination of egocentrism–the inability to consider perspectives other than one’s own [Piaget, 1977]. Piaget’s “Three Mountain Task” has became the standard for assessing level-2 perspective-taking [Piaget and Inhelder, 1957]. Commonly, 4-year-olds consistently fail on Three Mountain Task. This changes markedly as children enter the concrete operational stage. Children around age 6 could recognize perspectives different from their own. By ages 7–8, they could consistently and successfully identify the perspective of the other person [Piaget and Inhelder, 1969, Wimmer and Perner, 1983, Wellman et al., 2001, Liu et al., 2008].
The elimination of egocentrism via the acquisition of level-1 and level-2 perspective-taking is understood as the foundational instantiation of theory-of-mind in a child’s mind [Wellman, 2011, Schurz et al., 2015, Santiesteban et al., 2015, Rakoczy et al., 2018, Cole and Millett, 2019]. Building on these visually concrete perspective-taking abilities, children continue to develop other perspective-taking abilities, i.e. emotional perspective-taking, as their brains mature [Tomasello et al., 2005, Singer and Lamm, 2009]. Together, visual perspective-taking and other later developed faculties, i.e. altruism, empathy, mental simulation, set the foundation of intentionality understanding in our brains [Iacoboni, 2009, De Waal and Preston, 2017, Liu et al., 2017, Caviola et al., 2021, Ninomiya et al., 2020]. It is thus argued that perspective-taking grounds intentionality understanding in theory-of-mind [Wellman, 1992, Frith and Frith, 2006, Apperly, 2010]. In other words, humans develop to understand other people’s perspectives first before knowing what other people want.
In order to investigate intentionality understanding and perspective-taking abilities in Vision Language Models (VLMs), we have curated IntentBench and PerspectBench, which contains over 400 cognitive experiments grounded in real-world scenarios and classic cognitive tasks. We have aligned 37 models for our analysis. Unlike human intelligence, we find VLMs achieving high performance on intentionality understanding but lower performance on perspective-taking. Our results question the widely held notion in cognitive science literature that understanding intentionality requires perspective-taking in the relevant modality and cast light on the discrepancies and similarities between human and machine intelligence.
2 Methods
Figure 2: VLMs perform significantly better in intentionality understanding compared to perspective-taking. Paired samples t-test: $p=2.62\times 10^{-19}$, $t=17.651$
2.1 Dataset
We have curated IntentBench and PerspectBench which are composed of 441 cognitive experiments to investigate intentionality understanding and perspective-taking abilities in Vision Language Models. The project is part of our open-source agenda: Growing AI Like A Child [Li et al., 2024]. In PerspectBench, we have curated 32 multi-image or video format cognitive experiments and 209 single-image format cognitive experiments. In IntentBench, we have prepared 200 cognitive experiments, in which 100 experiments are based on real-world ambiguous social scenarios.
2.2 Cognitive Experiments
The Three Mountain Task first invented by Jean Piaget is widely used in developmental psychology laboratories as the gold standard for testing level-1 and level-2 perspective-taking abilities in children [Piaget and Inhelder, 1957, Johnson, 1975, Ford, 1979, Lamm et al., 2007]. In a standard Three Mountain Task assessment, a child is instructed to position oneself in front of a model featuring three mountains. These mountains vary in size and are distinguished by unique characteristics: one is covered in snow, another has a red cross at its peak, and the third is topped with a hut. The child is then asked to perform a complete 360-degree examination of the model. Subsequently, another individual is introduced and takes a different vantage point to observe the model. The child is presented with several photographs that showcase various viewpoints of the model and is tasked with identifying which photograph accurately represents what the other person sees. At around four years of age, children typically select the photograph that matches their own perspective. By six years old, they begin to acknowledge viewpoints that differ from their own, and by the ages of seven to eight, they are generally able to reliably identify the perspective of another individual [Piaget and Inhelder, 1969]. To test level-1 and level-2 perspective-taking in Vision Language Models, we develop the Three Mountain task into formats that are suitable for benchmarks with minimal confounding details while preserving real-life spatiality. In particular, we use groups of 3-4 commonly-seen elastic cans organized into different spatial patterns to mimic the mountain model. Like in the original task, we use a doll placed to face the organization from different angels as the object of perspective-taking.
Intentionality understanding is believed to be grounded by rudimentary theory-of-mind abilities like level-1 and level-2 perspective-taking [Wellman, 1992, Frith and Frith, 2006, Apperly, 2010]. In developmental psychology, a critical subset of intentionality understanding experiments involves tests of action understanding [Searle, 1979]. Several computational hypotheses are proposed on how one could understand other people’s actions: for example, action understanding could be computationally modeled as pure inference [Gweon and Schulz, 2011, Gweon, 2021], as mental action simulation [Brass et al., 2007], or as inverse planning [Baker et al., 2009]. Typically, cartoon stimuli built via physic simulation engine are used frequently in action understanding in developmental psychology [Liu et al., 2017, Shu et al., 2021]. These stimuli are incorporated into IntentBench. However, a common critique of cognitive psychology tasks is that they lack realism and have limited applicability to real-world situations [Gomez-Marin et al., 2014]. Drawing inspiration from COIG-CQIA and its Ruozhiba dataset, many real-world ambiguous scenarios are incorporated into IntentBench for explicitly testing intentionality understanding in ethological conditions [Bai et al., 2024]. All in all, PerspectBench and IntentBench are built as two batteries of aforementioned cognitive experiments adopted for the Vision Language Models (Figure 4, 5 and 6).
Figure 3: Differential performance changes in intentionality understanding and perspective-taking in Vision Language Models as their model sizes increase ($y=0.0599x+0.3925,r^{2}=0.2797$; perspective-taking: $y=-0.0057x+0.1437,r^{2}=0.0176$).
2.3 Model Selection and Experiment
We evaluate the intention inference capabilities of three categories of Visual Language Models (VLMs). To ensure a fair comparison, all VLMs are evaluated on their ability to reason over images and texts under a zero-shot generation task. Detailed documentation of inference methodology can be seen in the our agenda paper [Li et al., 2024].
A complete list of models is given in Table 1. Model size data are curated at the same time. The models are categorized as follows:
  1. 1.
    Open-source VLMs with Multi-Image Reasoning: Includes models with different sizes and other variants such as CogVLM Series [Hong et al., 2024], Qwen series [Bai et al., 2023, Wang et al., 2024], and Blip2 [Li et al., 2023], LLaVA-Next [Liu et al., 2024] , which are capable of reasoning over interleaved multiple images and texts.
  2. 2.
    Closed-source VLMs with Multi-Image Reasoning: Includes proprietary models such as GPT series ( GPT-4v, GPT-4-turbo, GPT-4o-mini), Gemini Series, and Claude Series [OpenAI, , Gemini, , claude, ]. These models also support reasoning across interleaved images and texts,
  3. 3.
    Open-source VLMs with single-Image Reasoning: Includes models designed to process a single image alongside continuous text, such as InstructBlip Series, LLaVA Series [Dai et al., 2023, Liu et al., 2023a, b]
Figure 4: Examples of Vision Language Models Assessed with IntentBench
Figure 5: Examples of Vision Language Models Assessed with PerspectBench
Prompt Example:
Question + Please answer with a number/ Yes or No/ the option’s letter from the given choices directly.
In total, we have processed 37 models for evaluation. All the model performances in intentionality understanding and perspective-taking together with human baseline performances are presented here (Figure 1). In order to analyze the reasoning abilities of VLMs, we ask the models to explain their answers after they have given the answers.
2.4 Human Baseline
We have recruited in total 22 participants, all of whom are college students proficient in English. Participants have been instructed to skip a question if the question is phrased ambiguously or is too complicated to answer in 90 seconds. This question has been marked as failed if the human participant does not provide an answer. For each question, at least 80% of the humans should get it right. Otherwise, we would modify the questions, and new annotators will complete the modified questions. The human baseline result of each question has been normalized by the number of participants who have effectively produced an answer for that question.
3 Results
Figure 6: More Examples of Vision Language Models Assessed with IntentBench
We have found a clear dissociation between the model performance on intentionality understanding and perspective-taking. Specifically, all assessed models perform much better on IntentBench as compared to PerspectBench (Figure 1 and 2). While some of the highest-performing models on IntentBench (e.g. GPT-4o) are able to reach near-human performance in intentionality understanding, they still perform poorly at perspective-taking (Figure 1). To quantify how Vision Language Models perform differently on intentionality understanding and perspective-taking, we perform a paired samples t-test on the accuracy scores of these models on the two datasets. The paired t-test has revealed a statistically significant difference between the models’ performance on the two tasks, with a t-statistic of $t=17.651$ and a p-value of $p=2.62\times 10^{-19}$ (Figure 2). The results have confirmed our central argument, that Vision Language Models perform differentially on intentionality understanding and perspective-taking, even though the two cognitive tasks have been regarded as two sides of one coin regarding human cognition in previous cognitive science literature.
Moreover, we have observed distinct performance changes in intentionality understanding and perspective-taking as Vision Language Models scale in size (Figure 3). To quantitatively analyze the impact of model size on performance, we have conducted a linear regression analysis of accuracy scores against model sizes across our 37 Vision Language Models: for intentionality understanding, the coefficient of determination is $R^{2}=0.2797$, with a slope of $0.0599$ and an offset of $0.3925$. For perspective-taking, the coefficient of determination is $R^{2}=0.0176$, with a slope of $-0.0057$ and an offset of $0.1437$. The slope of $0.0599$ for intentionality understanding indicates that as the model size increases, the ability of intentionality understanding also increase, which is consistent with the scaling law hypothesis. Intriguingly, for perspective-taking, the slope of $-0.0057$ actually indicates that as Vision Language Models scale in size, they actually don’t get better in taking others’ perspectives, which is a direct contradiction to the scaling law hypothesis.
We further ask, are abilities of intentionality understanding and perspective-taking correlated with each other in Vision Language Models? To investigate this, we have computed the Pearson and Spearman correlations between the two measures. The results show that there is essentially no relationship: the Pearson correlation coefficient is $0.0252$ with $p=0.882$ and the Spearman correlation coefficient is $0.0115$ with $p=0.946$, both indicating a lack of statistical significance. In other words, these two abilities appear to be largely independent within our assessment of Vision Language Models.
4 Discussions
Figure 7: Failure Modes of Vision Language Models Assessed with IntentBench
Our study has assessed VLMs’ ability to intentionality understanding and perspective-taking. Our results have shown that VLMs appear to be proficient in intentionality understanding while, surprisingly, performing significantly worse in perspective-taking. On the high end of theory-of-mind hierarchy, intentionality understanding is suggested to be supported by the ability to ascribe abstract mental representations of others by reasoning from the objects’ perspectives. Specifically for the tasks in IntentBench, such reasoning have to be operated on visual information, which supposedly requires perspective-taking. Our finding that the same Vision Language Model could perform competently on IntentBench but poorly on PerspectBench challenges this common proposition. Here, we provide two potential interpretations for this surprising finding, taking into consideration the results of recent investigations into both perspective-taking and intentionality understanding in the literature as well as ongoing debates.
The first interpretation is that VLMs may be able to infer the intentions of the people in the images without attempting to take their perspectives [Kilner, 2004]. Specifically, understanding the actions behind intentions has been shown to be highly contingent upon the contextual information among the scenarios in which the actions take place. If VLMs could successfully comprehend and register the contextual cues among the images, they may be able to exploit the contingency between what is presented in the surroundings and the actions depicted using associative learning without even attempting to take the positions of the actors [Kilner, 2004, Bianco et al., 2024]. However, exploitation of associative learning without understanding the comparative significance and causal links between the contextual priors could lead to mistakes in action understanding, especially when there are several possible and salient ways to understand the actions depicted. This is exemplified by VLMs’ failed performances on IntentBench (Figure 7).
The second interpretation concerns the possibility that Three Mountain Task requires certain abilities beyond what is demanded by intentionality understanding. Specifically, recent evidence indicates that Three Mountain Task requires the simultaneous confrontation of perspectives in visual reasoning, which is possibly more demanding than level-2 perspective-taking alone [Moll and Meltzoff, 2024]. In addition, it has also been suggested that cognitive processes independent of theory-of-mind might also contribute to the ability of visual perspective-taking [Michelon and Zacks, 2006, Kessler and Rutherford, 2010, Samson et al., 2010, Epley and Caruso, 2012, Surtees et al., 2013]. Under this view, level-1 and level-2 perspective-takings are in fact not cognitive prerequisites of intentionality understanding, but rather totally orthogonal processes. Our work empirically suggest that in Vision Language Models, intentionality understanding and perspective-taking recruit orthogonal computational processes. Following this view, our results contradict with the traditional cognitive science view in human studies.
5 Conclusion
Overall, our study is the first attempt at assessing VLMs’ performance on intentionality understanding and perspective-taking. More contrasting experiments are needed to be done in order to truly dissociate the contributions of different mechanisms of intentionality understanding and perspective-taking in humans and Vision Language Models. Our findings contradict with the commonly held belief in cognitive science that understanding intentionality necessitates perspective-taking within the relevant modality.
There might be two reasons for this contradiction: either Vision Language Models are fundamentally different forms of intelligence that do not capture the characteristics of human intelligence, or previous cognitive experiments have methodological shortcomings—such as failing to consider the directional growth of human intellect and the inability to conduct intervention studies due to the directional growth—that constrain our investigation on true human cognition, where Vision Language Models might otherwise serve as in silico models.
Acknowledgments and Disclosure of Funding
The project is not funded by any organizations.